### v2

[…]
The Heterogeneous Productivity Effects of Generative AIthanks: We would like to thank Yashdeep Dahiya for outstanding research assistance and Samantha Eyler-Driscoll for manuscript editing and proofreading. This paper supersedes a previous version titled “The Unintended Consequences of Censoring Digital Technology – Evidence from Italy’s ChatGPT Ban”, April 2023.
David Kreitmeir [^1] Department of Economics and SoDa Labs, Monash University (davidkreitmeir.github.io and david.kreitmeir1@monash.edu)
Paul A. Raschky [^2] Department of Economics and SoDa Labs, Monash University (praschky.github.io and paul.raschky@monash.edu)
We analyse the individual productivity effects of Italy’s ban on ChatGPT, a generative pretrained transformer chatbot. We compile data on the daily coding output quantity and quality of over 36,000 GitHub users in Italy and other European countries and combine these data with the sudden announcement of the ban in a difference-in-differences framework. Among the affected users in Italy, we find a short-term increase in output quantity and quality for less experienced users and a decrease in productivity on more routine tasks for experienced users.
JEL: D8, J24, O33
Keywords: artificial intelligence, productivity
1 Introduction
The public release of OpenAI’s ChatGPT provided near universal [^3] Countries where ChatGPT is not accessible include China, Eritrea, Iran, North Korea, Russia and Saudi Arabia, among others. access to generative artificial intelligence (AI) tools at no or very low cost. Its subsequent quick adoption [^4] According to OpenAI, by November 2023, ChatGPT was recording approximately 100 million weekly users. https://techcrunch.com/2023/11/06/openais-chatgpt-now-has-100-million-weekly-active-users/ broadened the discussion about the impact of generative AI on society and its potential to boost worker productivity by performing relatively complex tasks and producing (seemingly) novel output, all while requiring only minimal technological knowledge on the part of users. However, ChatGPT also has the tendency to produce wrong or faulty outputs (e.g., “hallucinations”) that, in the absence of expert knowledge, are difficult to detect and costly to rectify and might ultimately undermine the productivity of some workers (Dell’Acqua et al., 2023).
One of the focal points in the discussion on generative AI’s societal impact is its ability to create and generate new content and knowledge. Similarly to prior advances in AI, generative AI can enhance productivity by replacing more routine tasks (Brynjolfsson et al., 2023; Kanazawa et al., 2022; Noy and Zhang, 2023; Peng et al., 2023) or improving users’ decision accuracy (Kleinberg et al., 2017; Almog et al., 2024; Cho, 2023). The feature that sets this generation of AI apart from previous ones is that, with its access to the universe of online knowledge, generative AI combines domain-specific information with rules, lending it the ability to create new content and ultimately opening the possibility of extending the production possibility frontier beyond an individual’s current level of training or expertise.
However, the accuracy of current generative AI models’ performance in some tasks such as text summarization or generation, combined with its clarity and confidence of delivery, might create the illusion that it enhances productivity in other domains. Inaccurate, faulty or “hallucinated” output may not be immediately detected and could be used as an input in a knowledge worker’s subsequent production flow. For instance, Kabir et al. (2024) analysed ChatGPT’s response to 517 programming questions and found that 52% of its answers were incorrect and that the users presented with these answers overlooked these errors 39% of the time. Nevertheless, users still tend to prefer to use ChatGPT because of its comprehensive responses (Kabir et al., 2024) and the confident language of the responses (Li et al., 2023). The biggest online discussion and help forum for developers, stack overflow, has banned the use of LLM generated content on the forum because the rate of “getting correct answers” from these tools, it too low. [^5] https://meta.stackoverflow.com/questions/421831/policy-generative-ai-e-g-chatgpt-is-banned
For some tasks (e.g., content writing), the process of detecting faulty output or rectifying generative AI–driven errors might be quick, while for others (e.g., software development), the same process can be tedious and time consuming. [^6] While the motivation for the relatively early public release of ChatGPT and other large-language models (LLMs) was to improve their performance with the human-generated data collected from user interactions with the tools, their output quality remains noisy, and expert knowledge is often required to accurately judge this quality. Moreover, Chen et al. (2023) show that ChatGPT’s performance on a number of tasks, including generating code, actually declined from version 3.5 to version 4.0, calling into question whether its performance and accuracy will continuously improve over time. In addition, del Rio-Chanona et al. (2023) show that the widespread use of ChatGPT has led to a decline in usage of online help forums such as Stack Overflow, which in return will decrease the human-generated ground truth data that can be used to improve AI models. There are also wide differences in the accuracy of generative AI output, driven not only by the complexity of the underlying task but also the size and quality of the underlying training data. The tools can leverage a very large online text corpus to predict the next word in tasks such as creative writing and chatting, but the training data for software development and code creation are limited to a relatively small number of online forums (e.g., Stack Overflow), where the ground truth can be noisy. [^7] For example, for less routine, more complex and more niche questions, the answers provided on Stack Overflow are not necessarily correct or are just initial solution suggestions instead of working solutions. While these suggestions might have received upvotes, signaling to the LLM their “usefulness” for its training, in reality, the content of the answers and ChatGPT subsequent output might be only an untested and ultimately not functioning code routines.
In cases where the underlying task is more complex and the output requires accuracy to be ultimately useful (e.g. software development), relying on generative AI might prolong task completion and decrease workers’ output quality (Dell’Acqua et al., 2023). Such problems might be more acute among less experienced workers, who may have less domain knowledge and require more time to detect and correct errors. Less experienced workers might also be more prone to continue using the tool because the alternatives (e.g., acquiring the knowledge and skills themselves) appear to be even costlier.
In this paper, we use observational data to analyse the heterogeneous effects of ChatGPT on the output quantity and quality of experienced and less experienced software developers. In particular, we exploit Italy’s sudden announcement of a ChatGPT ban as a natural experiment to examine the ban’s short-run effects on GitHub users’ productivity. We find that the ban had no systematic effect on the overall output of more experienced developers and only some small negative effects on their output for more routine tasks (resolving issues and debugging). However, among less experienced users, the short-term lack of access to ChatGPT increased both the amount of output and its quality. For this group of users, the likelihood that we observe any output-related activity on GitHub is approximately 10% higher for the two business days following the ban. This effect size shrinks for the subsequent days. In the same vein, we find some tentative evidence that Internet users in Italy adapted fairly quickly to the legislation by increasing their use of virtual private networks (VPNs) and encrypted routing to circumvent the ban.
Finding high-frequency and consistently measured, observational data on the output of knowledge workers that is comparable across countries is challenging. We follow a the approach from a number of recent papers (e.g., McDermott and Hansen, 2021; Holub and Thies, 2023; Shen, 2023) that have already exploited the temporal granularity of GitHub activity data and used this data as a proxy for software developers’ productivity. When using this data to proxy productivity, we rely on a number of assumptions which we test in a subsequent robustness section. We show that the results are not driven by changes in working hours or the level of complexity of the task. Considering that the ban was implemented close to a major holiday in the treatment and control countries, we also show that the changes in behaviour are not driven by seasonal factors. Our results are also robust to the use of a set of alternative outcome variables and also analysing the effect at the user–repository–day level.
Our results present some first, nonexperimental empirical evidence on the effects of restricting access to generative AI on workers’ performance in more complex tasks. Importantly, we show that the effects of generative AI are heterogeneous by worker’s experience.
Our study complements the existing, largely experimental, literature on the effects of generative AI on worker productivity in less complex tasks (e.g., content writing, customer support), where generative AI output is less error-prone, by examining a setting with more complex tasks, where AI-generated output can be less accurate or more faulty. Existing work by Brynjolfsson et al. (2023) and Noy and Zhang (2023) has found mainly positive productivity effects of generative AI and stronger effects for less experienced workers in the contexts of customer support and content writing tasks. In contrast, our results suggest that, for more complex tasks (e.g., code development), [^8] Related work by Peng et al. (2023) and Chatterjee et al. (2024) has found positive effects of GitHub Copilot on both the productivity and job satisfaction of software developers. While GitHub Copilot is also an AI-based tool developed by GitHub and OpenAI, it is specifically a code completion tool. In contrast, ChatGPT is a chat-based (rather than auto-complete) tool and can be used to produce entirely new code segments/programs based on a human language prompt; in such cases, the accuracy of generative AI is highly variable (Kabir et al., 2024). generative AI does not necessarily boost the productivity of less experienced workers and can even decrease their output quantity and quality. Our results, therefore, confirm the findings of Dell’Acqua et al. (2023) in a controlled experiment environment that, for tasks beyond ChatGPT’s current capabilities, using ChatGPT increases the time a worker spends on a task. We complement their results by highlighting the differential effects by knowledge worker’s level of experience. Our finding of heterogeneous effects for more complex tasks also empirically complements the larger discussion in economics on technological change and inequality in the labour market (e.g., Acemoglu, 2002; Autor et al., 2003; Goldfarb and Tucker, 2019; Acemoglu and Restrepo, 2020), in particular the productivity and labour market effects of AI (e.g., Brynjolfsson et al., 2017; Agrawal et al., 2019; Acemoglu, 2021; Eloundou et al., 2023).
The paper is organised as follows: Section 2 provides some background on ChatGPT and the Italian ban on the technology in 2023. Section 3 describes the data. Section 4 presents empirical results on the ban’s effect on worker productivity, and Section 5 concludes.
2 ChatGPT and the Italian Ban
ChatGPT, an LLM created by US startup OpenAI, has been used by millions of people since it launched in November 2022. Trained on a vast corpus of text data from the Internet as it was in 2021, this large-scale AI language model uses a transformer-based neural network to process natural language. During the training process, the model learned to identify patterns and relationships between words, phrases, and sentences, enabling it to generate text.
ChatGPT is accessible via a public website (chatgpt.openai.com) or an application programming interface (API), and almost anyone [^9] Before Italy, countries including China, Russia and North Korea had already banned ChatGPT. can sign up for a free account. The interface is designed like a chat environment where the user writes “prompts” and ChatGPT answers. Interactions can range from casual chats and search-like queries to more complex exchanges such as creative writing of a text or creation of recipes based on prompts. ChatGPT can also write code in multiple programming languages on the basis of a simple prompt.
On April 1, 2023, the Italian data protection authority (Garante per la protezione dei dati personali) blocked use of the ChatGPT chatbot, citing privacy concerns, and announced an investigation into OpenAI’s compliance with the European Union’s General Data Protection Regulation (GDPR). In particular, the authority stated that there was no legal basis for the mass collection and storage of personal data to train the algorithms underlying the platform’s operation. [^10] Shiona McCallum, “ChatGPT banned in Italy over privacy concerns”, BBC 01/04/2023, https://www.bbc.com/news/technology-65139406 There are a number of reasons to expect that the ban did not have an effect on software developers. During the ban access to other Natural Language Processing (NLP) powered tools that are powered by Open AI, such as GitHub CoPilot, was not affected. In contrast to ChatGPT which was trained on the universe of available online text corpus, GitHub Copilot was specifically trained only on code-repositories and build to produce context-aware codes. It provides more accurate results and is also widely used by code-developers. It is important to note, that our study only estimates the effects of a ban of ChatGPT on developer’s output and not a ban on all LLM based coding support tools. Alternative generative AI tools such as Claude and Bard which were launched a few weeks prior to the ban were still accessible. There were also multiple ways for programmers to immediately circumvent the ban (i.e. VPN) and guides on how to do that where shared online. [^11] https://www.programmareinpython.it/blog/chatgpt-bloccato-in-italia-che-fare/
However, anecdotal evidence shows that shortly after the inception of the ban, Italian developers went online and complained about the disruption caused by restricting access to ChatGPT, “[…] a tool that has become an essential part of [their] daily routine” as software developers. [^12] Semero, Anto “ChatGPT Banned in Italy: Mamma Mia! What’s Going On?” https://medium.com/@antonellosemeraro/chatgpt-banned-in-italy-mamma-mia-whats-going-on-97c44284e331, April 2 2023. Concerns were raised that restricting access to ChatGPT during a time when the field of software development is increasingly fast-paced, poses a severe threat to the competitiveness Italian developers and businesses. [^13] https://www.programmareinpython.it/blog/chatgpt-bloccato-in-italia-che-fare/ Whether the ban was effective or not is ultimately an empirical question, which this paper aims to shed light on.
The ban was lifted in late April after OpenAI responded to the data protection authority’s privacy concerns. [^14] Shiona McCallum, “ChatGPT accessible again in Italy”, BBC 28/04/2023, https://www.bbc.com/news/technology-65431914
3 Data
GitHub Data
GitHub is the world’s largest online code hosting platform, used for storage of and joint work on coding projects (so-called repositories). [^15] The programming languages most commonly represented in GitHub repositories are Python (17.38%), Java (11.77%), Go (10%), JavaScript (9.95%), and C++ (9.66%). In comparison, R-related repositories account for only 0.074% of all pull requests on GitHub. https://madnight.github.io/githut All modifications to a GitHub repository are automatically timestamped and stored, and GitHub permits tracking of any iterations of specific files and lines of code. Every action taken by a team member is automatically recorded, with details about the kind and substance of the modification, the files and code lines affected, and the date the changes were performed. Anyone with access to a repository can examine and download the history of iterations and actions, and given GitHub’s history of developing open-source software, a significant portion of its repositories are not access restricted, meaning that the project activity information is available to everyone. Thus, public GitHub repositories provide a direct, real-time measure of labour activity for millions of software and code developers worldwide (McDermott and Hansen, 2021). [^16] GitHub data have been used in empirical research on software developers’ productivity during the onset of COVID-19 (Forsgren, 2021), the impact of COVID-19 on daily and weekly patterns of individual labour allocation (McDermott and Hansen, 2021), the effects of working from home on individual productivity (Shen, 2023), the effect of air pollution on individual output (Holub and Thies, 2023), and the relationship between social links and the likelihood of joining professional software development teams (Casalnuovo et al., 2015).
We access individual-level, real-time activity data from the GitHub Archive for GitHub users in Italy (treatment) and Austria, France, and Spain (control) in the week prior to and that immediately after the ChatGPT ban in Italy (March 27–April 9, 2023). [^17] We chose these three countries because all of them are part of the European Union and share a common land or sea border. In settings like ours, it is difficult to find objective criteria that help guiding the choice of comparable units. To provide further confidence that our results are not based on the choice of control countries, we conduct a leave-one-out analysis in Figure LABEL:fig:leave-one-out and show that the results are not driven by the composition of the control group. To account for the Easter break starting with Good Friday on 7 April—a public holiday in all four sample countries— we restrict the post-treatment period to Monday 3 April to Thursday 6 April and the corresponding pre-period to Monday 27 March to Thursday 30 March. [^18] For more details on the study’s time line and a graphical illustration please see A in the appendix. GitHub Archive is hosted on Google’s BigQuery warehouse system and contains all public event data of GitHub user, which is updated daily and can be accessed with a query on Google’s cloud infrastructure. GitHub user information such as the year of GitHub user account creation was downloaded with the GitHub GraphQL API. [^19] For more information on how we retrieve GitHub user location information, please see B.1 in the appendix. The two datasets are merged via the unique GitHub user login.
We use the individual-level action data to construct two sets of baseline outcome variables: The first group captures output quantity and quality and includes aggregate Output limited to “productive” actions, aggregate Output as defined by Shen (2023), aggregate Output as defined by Holub and Thies (2023), and the Pull request (PR) merge ratio quantifying how many of a user’s suggested code edits were accepted by the repository (project) owners. The second group gauges task choice and complexity and comprises PR opened, Avg. lines added per opened PR, Avg. lines added per merged PR, Easy issue closed, and Interactive activity. A detailed description of the construction and definition of all outcome variables is provided in appendix Table LABEL:tab:var-defs.
On a daily basis, GitHub actions are relatively rare events at the user level. Hence, we transform each count variable into a binary indicator that equals 1 if one of the actions in a category is recorded for the user on a given day, and 0 otherwise. [^20] The distribution of day–user-level counts of the main event variables during the sample period is presented in Figure B.1 in the appendix. Descriptive statistics at the user–day and user level are presented in appendix Table LABEL:tab:summary-user. To use GitHub actions as a measure for developers’ productivity, we need to ensure that the ban did not affect developers’ working hours. While it is not possible to have data on exact working hours for developers, we are still able to check if time spent on GitHub activity during the day has changed between the pre and post period. Using the timestamp data for each individual GitHub action, we are able to show that the distribution of GitHub activities across hours of the day did not change over the period (see Figure LABEL:fig:distribution-work-activity-Italy).
Package Repositories
We compile a list of packages hosted on GitHub for ten analytical programming languages: C, C++, Go, Java, JavaScript, Julia, Perl, Python, R, and Rust. We rely in the first instance on the community–curated “Awesome Lists” to locate GitHub repositories for “popular” packages in each language. In a second step, we scrape the information on all packages hosted on the official software repositories for Python (pypi), R (CRAN) and Julia (JuliaRegistries) to retrieve information on each package’s GitHub repository. We make use of the standardized GitHub URL structure to identify the owner and name of a package repository. [^21] The stylized URL for a package repository is https://GitHub.com/[account name hosting the repository]/[repository name] (for instance, https://github.com/numpy/numpy). To identify the GitHub user accounts other than the owner that contribute to a package repository, we use information on each individual GitHub user’s activity from January 2011 until March 2023. We restrict the list of GitHub event types to “productive” events to select primarily accounts that made at least one substantial contribution to a package repository. [^22] Our set of “productive” event types comprises PullRequestEvent, PullRequestReviewEvent, PullRequestReviewCommentEvent, PushEvent, and ReleaseEvent. Moreover, we winsorise the sample at the 1st and 99th percentiles to safeguard against outliers. [^23] Note that we exclude bot accounts from the list of contributors prior to winsorising. Our final list of package contributors and owners comprises $483,855$ unique GitHub user accounts, of which $5,916$ are part of our baseline sample.
4 Effect of the ChatGPT Ban on GitHub Output
To analyse the effect of the Italian ChatGPT ban on GitHub users’ output, we estimate variants of the following difference-in-difference (DID) specification:
$Y_{it}=$ $\beta D_{it}+\alpha_{i}+\lambda_{t}+\gamma_{dow}+\sigma_{d}\times t+\epsilon_{it},$ (1)
where $Y_{it}$ denotes the outcome variable, i.e. one of the user-specific output and task variables, and $D_{it}$ is the treatment indicator variable equaling 1 for Italian users in the first four work days post the ChatGPT ban. The parameters of interest is $\beta$. Time-invariant differences between users, including ability and experience, are captured by user fixed effects $\alpha_{i}$, while day (date) fixed effects $\lambda_{t}$ account for daily fluctuations in coding output across users. Additionally, we account for differences in working behaviour across week-days via day-of-the week $\gamma_{dow}$ and include control treatment and control group specific time trends to safeguard against differential time-trends in coding activity between Italian Github users and their European peers. [^24] We estimate alternative specifications of our baseline regression model. Including country-specific linear time-trends or excluding time-trends altogether leaves our baseline estimates quantitatively and qualitatively stable (Table LABEL:tab:did-wo-trend and Table LABEL:tab:did-ctry-trend in the appendix).
To check for potential pre-trends and investigate how the estimated effect evolves over time, rather than averaging over the whole window as in the generalized DID specification, we also estimate the following event-study specification:
$Y_{it}=$ $\sum_{\tau=-4}^{-2}\beta_{\tau}D_{it}^{\tau}+\sum_{\tau=0}^{3}\beta_{\tau}D_{it}^{\tau}+\alpha_{i}+\lambda_{t}+\gamma_{dow}+\sigma_{d}\times t+\epsilon_{it},$ (2)
where $D$ is a dummy variable equalling one for observations in the treatment group at event–day $\tau$ and zero otherwise; with $\tau=-1$ serving as the (excluded) reference period. In both specifications, we cluster standard errors at the user level.
4.1 Baseline Results
Table 1 presents our baseline DID model estimates for the coefficient of interest $\widehat{\beta}$. The upper panel reports the average treatment effect of the ChatGPT ban on output quantity and quality for four different samples (overall, less experienced, experienced and package contributors), while Panel B presents baseline estimates for task choice and complexity. We find that the ban has overall no significant effect on the output quantity and quality of Italian users or their choice of tasks with the notable exception of a negative and significant impact on their ability to close issues. These aggregate results, however, disguise important heterogeneity across users.
In Panel B, we provide suggestive evidence that the estimated effect on output quantity and quality of less experienced users is not result of a change in tasks. In particular, we do not see that the number of pull requests send for review (column 1) changes nor their complexity (column 2). Further, restricting the focus to lines added on pull requests that were actually merged (column 3), reveals a significant positive effect. While this increase could reflect less efficient code by Italian users post-ban, our estimates provide no support for the notion that the increase in the PR merge ratio is the result of decrease in merged PR task complexity. Similarly, we do not observe a switch to easier issues (column 4) nor to collaborative engagement instead of coding activity (column 5). [^25] In unreported results, we test if the estimated effect on easy issues is the result of restricting the attention to the resolution of those issues. If we consider any form of engagement with easy issues—i.e. commenting, opening or closing—we find that the probability of engaging with easy issues declines for less experienced users (significant at the 10% level), while no change is observed for more sophisticated users. For experienced users, we find null effects of the ban on task allocation and complexity, while the ban appears to have some effects on package contributors. In particular, the decline in the probability of opening a pull request and resolving easy issues might indicate that package contributors tackle less but harder tasks after the ban.
Figure 1 presents the estimated event-study coefficients $\beta_{\tau}$ from specification 2 for less experienced users. Importantly, a joint F-test of whether all coefficients prior to the ban are jointly zero cannot be rejected at conventional levels for any outcome, alleviating concerns about preexisting trends. Further, the event-study results reveal that the treatment effect for less experienced users peaked two days after the ban. Figure LABEL:fig:event-study-exp and LABEL:fig:event-study-pkg in the appendix present the corresponding event-study estimates for experienced users and package contributors, respectively. Results for this subset of users reveal that point estimates are rather volatile over time, particularly for package contributors, and reveal no clear pattern except for issue closed. Overall, our event-study results provide support for our main DID findings that the ban actually increased the productivity of less experienced coders, while more sophisticated users were largely unaffected.
4.2 Robustness Checks
We conduct a number of robustness checks and present the results in the appendix.
Alternative outcomes:
First, we consider a number of additional outcomes in Panel A of Table LABEL:tab:did-alt-out. We show that our results are largely robust to alternative measures for output quantity (Any Event and Commit), output quality (PR merged and PR merge ratio (Holub and Thies, 2023)) as well as task complexity (Avg. files edited per merged PR). In Panel B, we additionally show that our baseline findings are qualitatively stable at the intensive margin when we use continuous outcome variables instead of binary indicators.
Placebo tests:
Second, the overall changes in output quantity and quality could be driven by unobserved factors occurring at the same time as the introduction of the ban or general differences in Italian Github user activity relative to their European peers in the work week leading up to the Easter break. To address these concerns, we undertake two placebo exercises. In the first placebo test, we assume that the ban was implemented on the weekend prior to our actual pre-treatment period from Monday 27 to Thursday 30 March 2023. The estimated placebo treatment effects are presented in Table LABEL:tab:did-placebo-2023 and are relatively small in magnitude and vastly insignificant. Importantly, we don’t find the heterogeneous effects across users uncovered in our baseline analysis. Our second placebo treatment period spans the work week prior to Good Friday on 15 April 2022. Reassuringly, our placebo estimates (Table LABEL:tab:did-placebo-2022) are again exceedingly insignificant and do not follow a clear and comparable pattern to our baseline results.
Repository-level Analysis:
For a subset of users working on multiple repositories, we construct a new panel dataset at the user–repository–day level (Section LABEL:sec:repo-results in the appendix). [^26] Note that the set of public repositories is also restricted to those with at least one user from both the control and treatment group. This allows us to include repository (project) and user fixed effects. Exploiting only within user-repository variation projects we estimate the effect of the ban on users while holding the complexity of a project constant. The results in Table LABEL:tab:did-repo-main largely confirm the patterns from our baseline, user-level, analysis. For less experienced users, we still observe a positive and significant effect on output quantity and quality, while the likelihood of closing an issue for is still negatively affected by the ban. Interestingly, we now also find some suggestive evidence for the negative impact of the ban on experienced users’ output quantity relative to that of their European peers working on the same repository. Moreover, we find additional support for the negative effect of the ban on experienced users when conditioning on programming language and restrict our attention to highly complex official package repositories (Table LABEL:tab:did-repo-lang). [^27] Because of the limited number of observations for package repositories of other programming languages, our analysis is, unfortunately, constrained to official python packages from PyPi. In addition, we exclusively present results for output quantity and quality in LABEL:tab:did-repo-lang due to the limited variation in the data for most outcomes concerning task choice and complexity.
Additional Robustness Checks
We conduct a number of additional sensitivity checks. First, we conduct a ”Leave-One-Out” analysis to further alleviate concerns that our baseline findings are dependent on our selection of control group countries. In particular, we show in Figure LABEL:fig:leave-one-out that our results are robust to excluding users from each of the control group countries. Finally, we address concerns that our results could suffer from an overrejection of the null hypotheses as a result of reuse of the identifying exogenous variation for multiple outcomes (and subsamples). Table LABEL:tab:did-rwolf reports p-values (in brackets) that are corrected for multiple hypothesis testing following the procedure described in Romano and Wolf (2005a, b, 2016). [^28] In recent work, Heath et al. (2022) show that the employed Romano–Wolf correction procedure performs well in a multitude of settings and across different dimensions. Importantly, output quantity as defined in Shen (2023) and, in particular, output quality retain their significance even after correcting for multiple hypothesis testing.
4.3 Discussion
Data from public GitHub repositories include code and software projects from a variety of organisations and individuals. Some of these are open-source development projects (e.g., APIs) from private-sector companies, some are general open-source projects developed by a community of volunteers (and therefore are closer in character to public goods), and others are owned by research organisations or individual developers. Given data limitations, it is not possible to distinguish the type of project.
It is possible that some Italian users immediately used tools (e.g., VPNs) to circumvent the ban. Using data on Google searches for VPN services and usage data for TOR [^29] The TOR (The Onion Router) network is an open-source overlay network of thousands of network relays that conceals a user’s IP address. Unfortunately, we cannot access actual VPN usage data at daily level. (see appendix LABEL:sec:google-tor), we show a sudden jump in circumvention activity among Italian Internet users in the days after the ban. Despite the easy access to circumvention technology, many corporations and organisations actually prohibit the use of VPN and TOR tools on their devices and networks, implying that their use may be limited to mainly private devices and home networks. More importantly, we still find systematic effects on output despite this circumvention activity, and one can interpret our results as a lower bound. Another concern is that our finding of heterogeneity between less experienced and experienced users could be driven by the latter’s greater skill in circumventing the ban. However, the systematic effects of the ban on tasks related to closing issues and the negative output effect detected in repository-level analysis suggests that this is not the case.
There are a number of follow-up questions that we are unable to empirically analyse because of limitations in the data. First, while generative AI might disrupt the production flow of less experienced workers by providing faulty results, another possibility is that ChatGPT is simply a distraction. While it is not clear how this would explain the effect heterogeneity, more detailed data on the actual use of ChatGPT could help inform the design of workplace policies around generative AI. [^30] For example, existing generative AI tools such as GitHub Copilot are in general productivity enhancing because they are designed for specific tasks; Copilot, for example, only completes code and is not an open-ended chatbot. Second, more detailed data would also shed light on the question of why, after the initial increase in output and quality, we observe a decrease in the effect size for less experienced users in subsequent days. One explanation, in line with the conclusions of Kabir et al. (2024) and Li et al. (2023), could be that less experienced users still prefer to use ChatGPT as a support tool because it generates accessible and easy-to-use responses and because the costs of pursuing alternative solutions (e.g., acquiring the necessary coding skills) is relatively high. Finally, our study provides evidence on the productivity effects of (the ban on) generative AI in only the very short run because the ban was short-lived and circumventing it was relatively easy.
5 Conclusion
We present novel evidence of the short-term effects of generative AI (ChatGPT) on the productivity of knowledge workers using high-frequency, observational data from over 36,000 software developers in Italy and other European countries. We use the sudden ban on ChatGPT in Italy as a natural experiment and show that the access restriction distorted output quantity and quality. Our results not only present some first empirical evidence of the widespread adoption of ChatGPT in software and code development but also show that the productivity effects of ChatGPT (and restrictions on it) differ by experience level. Our findings have the following policy implications: For some, more complex tasks, generative AI can produce faulty and erroneous output that is difficult to detect, in particular for less experienced individuals. This calls for a more targeted use of the tool in both education and work. AI-based tools that harness the power of LLMs in a more controlled form, that generate a clearly defined output and that are not based on simple text prompts (e.g., GitHub Copilot) offer guard rails to ensure more domain-specific use. Our findings also indicate that even well-intended government-mandated blocking of digital technology (to protect privacy) can lead to short-term output disruptions and costs for society. Sudden bans can be easily circumvented with VPN tools, but these adjustment activities simultaneously distort production processes and negatively impact productivity in professions that rely on the banned technology. Thus, our research also implies that policymakers should consider the potential economic cost of digital technology bans before imposing them.
A: Output Quantity and Quality
Output Output (Shen, 2023) Output (Holub & Thies, 2023) Issue closed PR merge ratio
Overall Treated $\times$ Post 0.0062 0.0102 0.0090 -0.0045** 0.0045
(N = 290,864) (0.0076) (0.0075) (0.0080) (0.0022) (0.0030)
Dep. var. mean 0.2314 0.2283 0.2624 0.0142 0.0359
Less experienced Treated $\times$ Post 0.0193* 0.0216** 0.0203* -0.0017 0.0118***
(N = 149,680) (0.0107) (0.0107) (0.0111) (0.0023) (0.0035)
Dep. var. mean 0.2339 0.2316 0.2519 0.0089 0.0258
Experienced Treated $\times$ Post -0.0079 -0.0020 -0.0030 -0.0077* -0.0034
(N = 141,184) (0.0107) (0.0106) (0.0116) (0.0039) (0.0052)
Dep. var. mean 0.2287 0.2249 0.2734 0.0198 0.0467
Pkg. contributor Treated $\times$ Post 0.0181 0.0164 0.0270 -0.0151* 0.0113
(N = 47,328) (0.0205) (0.0200) (0.0215) (0.0084) (0.0104)
Dep. var. mean 0.2688 0.2644 0.3312 0.0267 0.0615
B: Task Choice and Complexity
PR opened Avg. lines added per opened PR Avg. lines added per merged PR Easy issue closed Interactive Activity
Overall Treated $\times$ Post $0$0.0039 0.0220 -0.0004 0.0006
(N = 290,864) (0.0034) (0.0148) (0.0135) (0.0004) (0.0043)
Dep. var. mean 0.0402 0.1581 0.1439 0.0007 0.0637
Less experienced Treated $\times$ Post 0.0040 0.0211 0.0456*** -0.0004 0.0037
(N = 149,680) (0.0040) (0.0186) (0.0164) (0.0004) (0.0045)
Dep. var. mean 0.0315 0.1358 0.1148 0.0005 0.0340
Experienced Treated $\times$ Post -0.0046 -0.0324 -0.0031 -0.0003 -0.0028
(N = 141,184) (0.0056) (0.0237) (0.0221) (0.0008) (0.0077)
Dep. var. mean 0.0494 0.1817 0.1749 0.0009 0.0953
Pkg. contributor Treated $\times$ Post -0.0221* -0.0933* 0.0226 -0.0021* 0.0127
(N = 47,328) (0.0117) (0.0477) (0.0422) (0.0011) (0.0158)
Dep. var. mean 0.0661 0.2356 0.2210 0.0014 0.1393
  • Notes:
  • All specifications include user–fixed effects, day-of-the-week–fixed effects, and a linear time trend for the control and treatment group. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Robust standard errors in parentheses are clustered on the user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
Figure 1: Less Experienced Users – Event-Study Estimates
Notes: Event-study estimates across outcomes for “less experienced” GitHub user accounts (created after or in 2017). The sample period spans March 27–30 (Pre) and April 3–6 (Post). All specifications include user, time, and day-of-the-week fixed effects. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). 95% (90%) confidence intervals for robust standard errors clustered at the user level are depicted in light (dark) grey.
Supplementary Online Appendix
[…]
Appendix A Timeline
Figure A.1 depicts the timing of the events. The ban was introduced on Friday March 31st 2023. In our baseline analysis, we consider the 4 business days prior to the ban, (MON – THU, March 27th – 30th) as the pre-period. To make the analysis, comparable, we ignore the two days following the introduction of the ban which were weekend days (SAT and SUN, April 1st – 2nd, in gray are excluded). The Post period is defined as the first 4 business days, after the introduction of the ban (MON – THU, April 3rd – 6th 2023). We further exclude Friday, 7th 2023, because this was Good Friday, a national holiday in all European countries under consideration in our sample.
Figure A.1: Time Line
PRE PeriodBan Intro/WE (Exc.)POST PeriodEaster (Exc.)MON 27/03THU 30/03FRI 31/03SUN 02/04MON 03/04THU 06/04
Appendix B Data Appendix
B.1 Github User Location
Figure B.1: GitHub User Profile
Notes: Profile last accessed on 07 May 2024 at https://github.com/0x0f0f0f.
We scrape, among others, the following key attributes at the user-level for a defined set of locations (location) and span of time-period (i.e. createdAt from 1 Jan 2009 until 11 Apr 2023) from the GitHub GraphQL API: [^31] For more details on the user-level attributes please refer to the official GitHub documentation: https://docs.github.com/en/graphql/reference/objects#user.
  • login: The user’s username (e.g. 0x0f0f0f).
  • location: The user’s location (e.g. Italy).
  • followers: The number of followers the user has (e.g. 165).
  • following: The numbers of users the user is following (e.g. 88).
  • repositories: A list of user’s (public) repositories (e.g. 44).
  • createdAt: The date and time the user’s account was created.
For more details on the scraping procedure (e.g. the set of locations) and the python implementation please refer to the public GitHub repository https://GitHub.com/sodalabsio/GitHub_scrape.
B.2 Definition and Construction of Outcome Variables
{ThreePartTable}
Table B.1: Variable Definitions
Variable Description
A – Output Quantity
Events Sum of all 15 GitHub event types, i.e., CommitCommentEvent, CreateEvent, DeleteEvent, ForkEvent, GollumEvent, IssuesEvent, IssueCommentEvent, MemberEvent, PublicEvent, PullRequestEvent, PullRequestReviewEvent, PullRequestReviewCommentEvent, PushEvent, ReleaseEvent, WatchEvent
Output # Commits [PushEvent\$.size] + # Issues closed [IssuesEvent\$.action == ’closed’]+ # Pull requests closed [PullRequestEvent\$.action == ’closed’] + # Releases [ReleaseEvent]
[…]
Issues closed # Issues closed [IssuesEvent\$.action == ’closed’]
Commits # Commits [PushEvent\$.size]
B – Output Quality
Pull Requests (PRs) merged # Closed pull requests that were merged [PullRequestEvent\$.pull_request.merged == ’true’ & PullRequestEvent\$.action == ’closed’]
PR merge ratio ( # Closed pull requests that were merged [PullRequestEvent\$.pull_request.merged == ’true’ & PullRequestEvent\$.action == ’closed’]) / (# Pull Requests closed [PullRequestEvent\$.action == ’closed’])
PR merge ratio (Holub and Thies, 2023) ( # Opened pull requests that were merged [PullRequest\$.pull_request.merged == ’true’ & PullRequestEvent\$.action == ’closed’ & PullRequestEvent\$.action == ’opened’]) / (# Pull Requests opened [PullRequestEvent\$.action == ’opened’])
C – Task Choice and Complexity
Pull Requests (PRs) opened # Pull Requests opened [PullRequestEvent\$.action == ’opened’]
Avg. files edited per merged PR Average # files edited per merged pull request [AVG(PullRequestEvent\$.pull_request.changed_files) IF PullRequestEvent\$.pull_request.merged == ’true’ & PullRequestEvent\$.action == ’closed’]
Avg. Lines added per merged PR Average # lines added per merged pull request [AVG(PullRequestEvent\$.pull_request.additions) IF PullRequestEvent\$.pull_request.merged == ’true’ & PullRequestEvent\$.action == ’closed’]
Avg. Lines added per opened PR Average # lines added per opened pull request [AVG(PullRequestEvent\$.pull_request.additions) IF PullRequestEvent\$.action == ’opened’]
Easy issue closed # Easy issues closed [IssuesEvent\$.action == ’closed’]
Interactive Activity # Comments on issues [IssueCommentEvent]+ # Comments on pull requests [PullRequestReviewCommentEvent] + # Comments on commits [CommitCommentEvent]
D – User Activity
“Work” hours Time difference (in hours) between first and last activity (any or “prodctive”)]
First activity Clock Hour of first activity (any or “prodctive”)
Last activity Clock Hour of last activity (any or “prodctive”)
Unusual activity (10th %tile) Dummy = 1 if activity (any or “prodctive”) before 6am (10th percentile of First activity) or after 9pm (90th percentile of Last activity)
Unusual activity (25th %tile) Dummy = 1 if activity (any or “prodctive”) before 8am (25th percentile of First activity) or after 7pm (75th percentile of Last activity)
[…]
Appendix C GitHub User-Level Data
C.1 Descriptive Statistics
Overall Less Experienced Experienced Pkg. contributor
(N = 36,358) (N = 18,710) (N = 17,648) (N = 5,916)
Mean SD Mean SD Mean SD Mean SD
A - User-Day–Level (N = 290,864)
     Output 0.2314 0.4217 0.2339 0.4233 0.2287 0.4200 0.2688 0.4433
     Output (Shen, 2023) 0.2283 0.4198 0.2316 0.4218 0.2249 0.4175 0.2644 0.4410
     Output (Holub and Thies, 2023) 0.2624 0.4399 0.2519 0.4341 0.2734 0.4457 0.3312 0.4706
     Issue closed 0.0142 0.1183 0.0089 0.0938 0.0198 0.1394 0.0267 0.1611
     PR merge ratio 0.0359 0.1846 0.0258 0.1572 0.0467 0.2092 0.0615 0.2381
     PR opened 0.0402 0.1964 0.0315 0.1746 0.0494 0.2168 0.0661 0.2484
     Avg. lines added per opened PR 0.1581 0.9003 0.1358 0.8648 0.1817 0.9359 0.2356 1.0474
     Avg. lines added per merged PR 0.1439 0.8563 0.1148 0.7928 0.1749 0.9179 0.2210 1.0124
     Easy issue closed 0.0007 0.0263 0.0005 0.0219 0.0009 0.0302 0.0014 0.0367
     Interactive Activity 0.0637 0.2442 0.0340 0.1811 0.0953 0.2936 0.1393 0.3463
B - User–Level (N = 36,358)
Mean SD Min Median Max
     User creation year 2016.55 3.77 2009 2017 2023
     Experienced 0.49 0.50 0 0 1
     Pkg. contributions 49.81 447.12 0 0 19638
     Pkg. owner 0.05 0.22 0 0 1
     Followers 29.49 203.22 0 6 17421
     Following 19.67 185.41 0 5 28300
     Repositories 29.90 54.53 0 17 3900
     Total events 11.93 18.87 1 5 140
  • Notes:
  • Panel A presents descriptive statistics for the baseline sample period Pre 27-30.03 – Post 03-06.04. The “Less experienced” sample includes all GitHub user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of unique GitHub user accounts for the entire baseline sample (“Overall”) and each of the subsamples is presented in parentheses below. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Panel B provides information on the individual characteristics of all GitHub user accounts in the baseline sample.
Notes: Daily counts of each action type at the user level for the sample period of March 27–30 (Pre) – April 3–6 (Post) are presented. Counts above 20 are binned and labelled $20+$.
Notes: The distribution of (any) activity by GitHub users across hours of the day is presented separately for the four days pre- and post-ban.
Notes: The distribution of (any) activity by GitHub users across hours of the day is presented separately for the four days pre- and post-ban.
C.2 Additional Results
(1) (2) (3) (4) (5)
A: Output Quantity and Quality
Output Output (Shen, 2023) Output (Holub & Thies, 2023 Issue closed PR merge ratio
Overall Treated $\times$ Post 0.0116*** 0.0128*** 0.0147*** -0.0005 0.0056***
(N = 290,864) (0.0043) (0.0042) (0.0045) (0.0011) (0.0015)
Dep. var. mean 0.2314 0.2283 0.2624 0.0142 0.0359
Less experienced Treated $\times$ Post 0.0177*** 0.0177*** 0.0186*** 0.0027** 0.0095***
(N = 149,680) (0.0061) (0.0061) (0.0062) (0.0013) (0.0018)
Dep. var. mean 0.2339 0.2316 0.2519 0.0089 0.0258
Experienced Treated $\times$ Post 0.0052 0.0079 0.0108* -0.0043** 0.0013
(N = 141,184) (0.0059) (0.0059) (0.0064) (0.0020) (0.0025)
Dep. var. mean 0.2287 0.2249 0.2734 0.0198 0.0467
Pkg. contributor Treated $\times$ Post 0.0148 0.0150 0.0194* -0.0095** -0.0001
(N = 47,328) (0.0109) (0.0108) (0.0115) (0.0040) (0.0049)
Dep. var. mean 0.2688 0.2644 0.3312 0.0267 0.0615
B: Task Choice and Complexity
PR opened Avg. lines added per opened PR Avg. lines added per merged PR Easy issue closed Interactive Activity
Overall Treated $\times$ Post 0.0035** 0.0180** 0.0249*** 0.0002 0.0011
(0.0017) (0.0078) (0.0070) (0.0002) (0.0022)
(N = 290,864) Dep. var. mean 0.0402 0.1581 0.1439 0.0007 0.0637
Less experienced Treated $\times$ Post 0.0056*** 0.0322*** 0.0402*** 0.0002 0.0016
(N = 149,680) (0.0021) (0.0100) (0.0087) (0.0002) (0.0024)
Dep. var. mean 0.0315 0.1358 0.1148 0.0005 0.0340
Experienced Treated $\times$ Post 0.0012 0.0025 0.0079 0.0001 0.0004
(N = 141,184) (0.0029) (0.0121) (0.0112) (0.0004) (0.0039)
Dep. var. mean 0.0494 0.1817 0.1749 0.0009 0.0953
Pkg. contributor Treated $\times$ Post -0.0083 -0.0269 -0.0006 -0.0010* 0.0005
(N = 47,328) (0.0057) (0.0240) (0.0219) (0.0006) (0.0081)
Dep. var. mean 0.0661 0.2356 0.2210 0.0014 0.1393
  • Notes:
  • All specifications include user–fixed effects and day-of-the-week–fixed effects. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Robust standard errors in parentheses are clustered on the user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
(1) (2) (3) (4) (5)
A: Output Quantity and Quality
Output Output (Shen, 2023) Output (Holub & Thies, 2023 Issue closed PR merge ratio
Overall Treated $\times$ Post 0.0062 0.0102 0.0090 -0.0045** 0.0045
(N = 290,864) (0.0076) (0.0075) (0.0080) (0.0022) (0.0030)
Dep. var. mean 0.2314 0.2283 0.2624 0.0142 0.0359
Less experienced Treated $\times$ Post 0.0193* 0.0216** 0.0203* -0.0017 0.0118***
(N = 149,680) (0.0107) (0.0107) (0.0111) (0.0023) (0.0035)
Dep. var. mean 0.2339 0.2316 0.2519 0.0089 0.0258
Experienced Treated $\times$ Post -0.0079 -0.0020 -0.0030 -0.0077* -0.0034
(N = 141,184) (0.0107) (0.0106) (0.0116) (0.0039) (0.0052)
Dep. var. mean 0.2287 0.2249 0.2734 0.0198 0.0467
Pkg. contributor Treated $\times$ Post 0.0181 0.0164 0.0270 -0.0151* 0.0113
(N = 47,328) (0.0205) (0.0200) (0.0215) (0.0084) (0.0104)
Dep. var. mean 0.2688 0.2644 0.3312 0.0267 0.0615
B: Task Choice and Complexity
PR opened Avg. lines added per opened PR Avg. lines added per merged PR Easy issue closed Interactive Activity
Overall Treated $\times$ Post $0$0.0039 0.0220 -0.0004 0.0006
(N = 290,864) (0.0034) (0.0148) (0.0135) (0.0004) (0.0043)
Dep. var. mean 0.0402 0.1581 0.1439 0.0007 0.0637
Less experienced Treated $\times$ Post 0.0040 0.0211 0.0456*** -0.0004 0.0037
(N = 149,680) (0.0040) (0.0186) (0.0164) (0.0004) (0.0045)
Dep. var. mean 0.0315 0.1358 0.1148 0.0005 0.0340
Experienced Treated $\times$ Post -0.0046 -0.0324 -0.0031 -0.0003 -0.0028
(N = 141,184) (0.0056) (0.0237) (0.0221) (0.0008) (0.0077)
Dep. var. mean 0.0494 0.1817 0.1749 0.0009 0.0953
Pkg. contributor Treated $\times$ Post -0.0221* -0.0933* 0.0226 -0.0021* 0.0127
(N = 47,328) (0.0117) (0.0477) (0.0422) (0.0011) (0.0158)
Dep. var. mean 0.0661 0.2356 0.2210 0.0014 0.1393
  • Notes:
  • All specifications include user–fixed effects, day-of-the-week–fixed effects and country-specific linear time trends. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Robust standard errors in parentheses are clustered on the user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
Notes: Event-study estimates across outcomes for “experienced” GitHub user accounts (created before 2017). The sample period spans March 27–30 (Pre) and April 3–6 (Post). All specifications include user, time, and day-of-the-week fixed effects. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). 95% (90%) confidence intervals for robust standard errors clustered at the user level are depicted in light (dark) grey.
Notes: Event-study estimates across outcomes for “pkg. contributor” GitHub user accounts (owner and/or contributor to a programming package repository). The sample period spans March 27–30 (Pre) and April 3–6 (Post). All specifications include user, time, and day-of-the-week fixed effects. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). 95% (90%) confidence intervals for robust standard errors clustered at the user level are depicted in light (dark) grey.
(1) (2) (3) (4) (5)
A: Additional Outcomes
Any Event Commit PR merged PR merge ratio (Holub & Thies, 2023) Avg. files edited per merged PR
Overall Treated $\times$ Post 0.0091 0.0084 0.0047 0.0017 0.0119**
(N = 290,864) (0.0090) (0.0075) (0.0031) (0.0018) (0.0060)
Dep. var. mean 0.3358 0.2215 0.0371 0.0130 0.0630
Less experienced Treated $\times$ Post 0.0282** 0.0220** 0.0118*** 0.0037 0.0224***
(N = 149,680) (0.0124) (0.0106) (0.0035) (0.0024) (0.0074)
Dep. var. mean 0.3191 0.2272 0.0265 0.0122 0.0498
Experienced Treated $\times$ Post -0.0116 -0.0061 -0.0029 -0.0004 0.0005
(N = 141,184) (0.0133) (0.0105) (0.0053) (0.0028) (0.0098)
Dep. var. mean 0.3534 0.2155 0.0483 0.0138 0.0769
Pkg. contributor Treated $\times$ Post 0.0311 0.0242 0.0107 -0.0024 0.0277
(N = 47,328) (0.0233) (0.0199) (0.0107) (0.0050) (0.0195)
Dep. var. mean 0.4088 0.2521 0.0635 0.0171 0.0971
B: Intensive Margin
Outputs Outputs (Shen, 2023) Outputs (Holub & Thies, 2023) PRs merged PRs opened
Overall Treated $\times$ Post 0.0122 0.0164 0.0180 0.0053* 0.0010
(N = 290,864) (0.0145) (0.0146) (0.0149) (0.0032) (0.0030)
Dep. var. mean 0.3706 0.3727 0.4215 0.0347 0.0345
Less experienced Treated $\times$ Post 0.0407** 0.0464** 0.0478** 0.0144*** 0.0041
(N = 149,680) (0.0190) (0.0191) (0.0193) (0.0036) (0.0035)
Dep. var. mean 0.3538 0.3545 0.3830 0.0247 0.0276
Experienced Treated $\times$ Post -0.0191 -0.0164 -0.0144 -0.0045 -0.0025
(N = 141,184) (0.0222) (0.0224) (0.0231) (0.0053) (0.0049)
Dep. var. mean 0.3884 0.3919 0.4624 0.0454 0.0418
Pkg. contributor Treated $\times$ Post -0.0002 0.0033 0.0313 0.0108 -0.0166*
(N = 47,328) (0.0445) (0.0445) (0.0454) (0.0108) (0.0101)
Dep. var. mean 0.4802 0.4852 0.5865 0.0589 0.0556
  • Notes:
  • All specifications include user–fixed effects, day-of-the-week–fixed effects, and a linear time trend for control and treatment group. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. files edited per PR and all intensive margin outcomes. Robust standard errors in parentheses are clustered on the user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
(1) (2) (3) (4) (5)
A: Output Quantity and Quality
Output Output (Shen, 2023) Output (Holub & Thies, 2023) Issue closed PR merge ratio
Overall Treated $\times$ Post -0.0010 -0.0014 -0.0052 -0.0016 -0.0070**
(N = 290,864) (0.0070) (0.0069) (0.0072) (0.0022) (0.0029)
Dep. var. mean 0.2164 0.2138 0.2444 0.0136 0.0347
Less experienced Treated $\times$ Post -0.0031 -0.0028 -0.0038 -0.0021 -0.0061*
(N = 149,680) (0.0097) (0.0096) (0.0099) (0.0023) (0.0033)
Dep. var. mean 0.2163 0.2144 0.2323 0.0083 0.0249
Experienced Treated $\times$ Post 0.0004 -0.0010 -0.0078 -0.0010 -0.0076
(N = 141,184) (0.0100) (0.0099) (0.0106) (0.0038) (0.0049)
Dep. var. mean 0.2164 0.2131 0.2573 0.0192 0.0452
Pkg. contributor Treated $\times$ Post 0.0096 0.0094 $0$0.0012 0.0020
(N = 47,328) (0.0189) (0.0185) (0.0198) (0.0082) (0.0099)
Dep. var. mean 0.2627 0.2582 0.3208 0.0271 0.0604
B: Task Choice and Complexity
PR opened Avg. lines added per opened PR Avg. lines added per merged PR Easy issue closed Interactive Activity
Overall Treated $\times$ Post -0.0043 -0.0226 -0.0206 -0.0006 -0.0044
(N = 290,864) (0.0032) (0.0142) (0.0131) (0.0005) (0.0039)
Dep. var. mean 0.0381 0.1509 0.1383 0.0006 0.0606
Less experienced Treated $\times$ Post -0.0069* -0.0246 -0.0308* -0.0006 -0.0027
(N = 149,680) (0.0037) (0.0178) (0.0157) (0.0005) (0.0042)
Dep. var. mean 0.0300 0.1285 0.1087 0.0004 0.0315
Experienced Treated $\times$ Post -0.0015 -0.0213 -0.0079 -0.0007 -0.0062
(N = 141,184) (0.0053) (0.0228) (0.0217) (0.0009) (0.0069)
Dep. var. mean 0.0467 0.1747 0.1696 0.0009 0.0913
Pkg. contributor Treated $\times$ Post -0.0018 -0.0397 0.0249 -0.0021 0.0002
(N = 47,328) (0.0107) (0.0438) (0.0427) (0.0018) (0.0149)
Dep. var. mean 0.0656 0.2366 0.2193 0.0012 0.1376
  • Notes:
  • The placebo post-treatment period ranges from Mon 27 March 2023 to Thu 30 March 2023; the corresponding pre-treatment period is from Mon 20 March 2023 to Thu 23 March 2023. All specifications include user–fixed effects, day-of-the-week–fixed effects, and a linear time trend for control and treatment group. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Robust standard errors in parentheses are clustered on the user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
(1) (2) (3) (4) (5)
A: Output Quantity and Quality
Output Output (Shen, 2023) Output (Holub & Thies, 2023 Issue closed PR merge ratio
Overall Treated $\times$ Post 0.0109 0.0100 0.0083 -0.0008 0.0034
(N = 145,480) (0.0110) (0.0109) (0.0115) (0.0037) (0.0049)
Dep. var. mean 0.2492 0.2455 0.2872 0.0196 0.0424
Less experienced Treated $\times$ Post 0.0159 0.0152 0.0152 -0.0051 0.0054
(N = 58,400) (0.0174) (0.0174) (0.0179) (0.0045) (0.0061)
Dep. var. mean 0.2448 0.2424 0.2669 0.0114 0.0274
Experienced Treated $\times$ Post 0.0079 0.0069 0.0037 0.0024 0.0023
(N = 87,080) (0.0142) (0.0141) (0.0149) (0.0054) (0.0071)
Dep. var. mean 0.2522 0.2475 0.3008 0.0252 0.0524
Pkg. contributor Treated $\times$ Post 0.0326 0.0261 0.0225 0.0077 -0.0015
(N = 35,680) (0.0243) (0.0240) (0.0252) (0.0103) (0.0127)
Dep. var. mean 0.2853 0.2800 0.3456 0.0311 0.0609
B: Task Choice and Complexity
PR opened Avg. lines added per opened PR Avg. lines added per merged PR Easy issue closed Interactive Activity
Overall Treated $\times$ Post 0.0021 0.0240 0.0413* $0$0.0071
(N = 145,480) (0.0050) (0.0232) (0.0214) (0.0005) (0.0066)
Dep. var. mean 0.0446 0.1718 0.1593 0.0008 0.0845
Less experienced Treated $\times$ Post 0.0028 0.0171 0.0354 -0.0003 -0.0037
(N = 58,400) (0.0067) (0.0338) (0.0271) (0.0005) (0.0081)
Dep. var. mean 0.0330 0.1382 0.1117 0.0006 0.0441
Experienced Treated $\times$ Post 0.0018 0.0297 0.0472 0.0002 -0.0095
(N = 87,080) (0.0070) (0.0315) (0.0310) (0.0008) (0.0096)
Dep. var. mean 0.0523 0.1943 0.1913 0.0010 0.1116
Pkg. contributor Treated $\times$ Post 0.0006 0.0170 0.0462 -0.0010 0.0037
(N = 35,680) (0.0128) (0.0572) (0.0493) (0.0012) (0.0178)
Dep. var. mean 0.0665 0.2376 0.2159 0.0014 0.1443
  • Notes:
  • The placebo post-treatment period ranges from Mon 11 April 2022 to Thu 14 April 2022; the corresponding pre-treatment period is from Mon 4 April 2022 to Thu 7 April 2022. All specifications include user–fixed effects, day-of-the-week–fixed effects, and a linear time trend for control and treatment group. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Robust standard errors in parentheses are clustered on the user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
(1) (2) (3) (4) (5)
A: Any Activity
“Work” hours First activity Last activity Unusual activity (10th %tile) Unusual activity (25th %tile)
Overall Treated $\times$ Post 0.0356 -0.2048 -0.1933 0.0032 0.0053
(N = 290,864) (0.0497) (0.1777) (0.1784) (0.0038) (0.0063)
Dep. var. mean 0.9620 11.3925 14.5319 0.0494 0.1292
Less experienced Treated $\times$ Post 0.1344** -0.2833 -0.0351 0.0083 0.0130
(N = 149,680) (0.0643) (0.2543) (0.2506) (0.0052) (0.0084)
Dep. var. mean 0.8479 11.7579 14.6761 0.0477 0.1196
Experienced Treated $\times$ Post -0.0742 -0.1427 -0.3592 -0.0026 -0.0029
(N = 141,184) (0.0772) (0.2489) (0.2543) (0.0055) (0.0094)
Dep. var. mean 1.0829 11.0537 14.3982 0.0512 0.1395
Pkg. contributor Treated $\times$ Post 0.1684 -0.0484 -0.1477 0.0069 0.0226
(N = 47,328) (0.1535) (0.3616) (0.3987) (0.0107) (0.0180)
Dep. var. mean 1.4032 10.8529 14.4870 0.0584 0.1652
B: Productive Activities
“Work” hours First activity Last activity Unusual activity (10th %tile) Unusual activity (25th %tile)
Overall Treated $\times$ Post 0.0346 0.0159 0.0083 -0.0004 -0.0010
(N = 290,864) (0.0428) (0.2032) (0.2017) (0.0032) (0.0053)
Dep. var. mean 0.7395 11.4108 14.6531 0.0374 0.0955
Less experienced Treated $\times$ Post $0*$2090 -0.0059 0.0031 0.0082
(N = 149,680) (0.0567) (0.2848) (0.2782) (0.0045) (0.0073)
Dep. var. mean 0.6884 11.7987 14.7975 0.0383 0.0946
Experienced Treated $\times$ Post -0.0591 0.2347 0.0107 -0.0043 -0.0110
(N = 141,184) (0.0651) (0.2906) (0.2937) (0.0045) (0.0077)
Dep. var. mean 0.7937 11.0049 14.5019 0.0365 0.0965
Pkg. contributor Treated $\times$ Post 0.0822 0.0431 -0.2626 -0.0041 0.0075
(N = 47,328) (0.1296) (0.4233) (0.4664) (0.0086) (0.0150)
Dep. var. mean 0.9935 10.8366 14.4380 0.0416 0.1147
  • Notes:
  • All specifications include user–fixed effects, day-of-the-week–fixed effects, and a linear time trend for control and treatment group. “Productive activities” comprise: PullRequestEvent, PullRequestReviewEvent, PullRequestReviewCommentEvent, PushEvent, ReleaseEvent, CreateEvent, IssueEvent. The number of observations is depicted in parentheses after each sample definition. For first and last activity only users with at least one activity on the day are included; the number of observations for the “overall”, “less experienced”, “experienced”, and “pkg. contributor” sample in Panel A, respectively, Panel B are 43,868, 40,672, 84,540, 17,802 and 30,717, 32,135, 62,852, 12,663. Robust standard errors in parentheses are clustered on the user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
Notes: Treatment effect estimates for the baseline DID specification are presented when users from one of the control group countries—i.e. Austria, France or Spain—are consecutively dropped from the sample. Point estimates and 95% (90%) cluster-robust confidence intervals are depicted in light (dark) grey for “Leave-One-Out” estimates and juxtaposed to the “Original” DID estimates (c. Table 1) in light (dark) red.
(1) (2) (3) (4) (5)
A: Output Quantity and Quality
Output Output (Shen, 2023) Output (Holub & Thies, 2023 Issue closed PR merge ratio
Overall (N = 290,864) Treated $\times$ Post 0.0062 0.0102 0.0090 -0.0045** 0.0045
[0.960] [0.602] [0.823] [0.083] [0.426]
Dep. var. mean 0.2314 0.2283 0.2624 0.0142 0.0359
Less experienced (N = 149,680) Treated $\times$ Post 0.0193* 0.0216** 0.0203* -0.0017 0.0118***
[0.179] [0.085] [0.157] [0.960] [0.000]
Dep. var. mean 0.2339 0.2316 0.2519 0.0089 0.0258
Experienced (N = 141,184) Treated $\times$ Post -0.0079 -0.0020 -0.0030 -0.0077* -0.0034
[0.960] [0.999] [0.998] [0.110] [0.971]
Dep. var. mean 0.2287 0.2249 0.2734 0.0198 0.0467
Pkg. contributor (N = 47,328) Treated $\times$ Post 0.0181 0.0164 0.0270 -0.0151* 0.0113
[0.945] [0.960] [0.710] [0.179] [0.835]
Dep. var. mean 0.2688 0.2644 0.3312 0.0267 0.0615
B: Task Choice and Complexity
PR opened Avg. lines added per opened PR Avg. lines added per merged PR Easy issue closed Interactive Activity
Overall (N = 290,864) Treated $\times$ Post $0$0.0039 0.0220 -0.0004 0.0006
[0.999] [0.998] [0.311] [0.956] [0.999]
Dep. var. mean 0.0402 0.1581 0.1439 0.0007 0.0637
Less experienced (N = 149,680) Treated $\times$ Post 0.0040 0.0211 0.0456*** -0.0004 0.0037
[0.898] [0.823] [0.004] [0.823] [0.960]
Dep. var. mean 0.0315 0.1358 0.1148 0.0005 0.0340
Experienced (N = 141,184) Treated $\times$ Post -0.0046 -0.0324 -0.0031 -0.0003 -0.0028
[0.960] [0.581] [0.999] [0.996] [0.996]
Dep. var. mean 0.0494 0.1817 0.1749 0.0009 0.0953
Pkg. contributor (N = 47,328) Treated $\times$ Post -0.0221* -0.0933* 0.0226 -0.0021* 0.0127
[0.132] [0.109] [0.989] [0.126] [0.960]
Dep. var. mean 0.0661 0.2356 0.2210 0.0014 0.1393
  • Notes:
[…]
Appendix D GitHub Repository–User-Level Data
Overall Less experienced Experienced Pkg. contributor
(N = 4,627 $\times$ 11,938) (N = 3,315 $\times$ 4,566) (N = 4,294 $\times$ 7,372) (N = 3,006 $\times$ 2,902)
Mean SD Mean SD Mean SD Mean SD
A - Repo-User-Day–Level (N = 239,912)
     Output 0.0116 0.1069 0.0080 0.0889 0.0136 0.1158 0.0171 0.1296
     Output (Shen, 2023) 0.0108 0.1036 0.0075 0.0860 0.0128 0.1122 0.0163 0.1264
     Output (Holub and Thies, 2023) 0.0279 0.1647 0.0192 0.1373 0.0327 0.1780 0.0405 0.1971
     Issue closed 0.0024 0.0492 0.0014 0.0376 0.0030 0.0546 0.0034 0.0580
     PR merge ratio 0.0030 0.0542 0.0013 0.0358 0.0039 0.0621 0.0047 0.0679
     PR opened 0.0054 0.0734 0.0041 0.0641 0.0061 0.0780 0.0080 0.0890
     Avg. lines added per opened PR 0.0170 0.2762 0.0134 0.2494 0.0190 0.2902 0.0238 0.3190
     Avg. lines added per merged PR 0.0104 0.2167 0.0047 0.1483 0.0135 0.2469 0.0153 0.2541
     Easy issue closed 0.0001 0.0094 0.0000 0.0034 0.0001 0.0114 0.0001 0.0089
     Interactive Activity 0.0193 0.1377 0.0127 0.1120 0.0230 0.1500 0.0285 0.1663
B - User–Level (N = 11,938)
Mean SD Min Median Max
     User creation year 2015.37 3.51 2009 2015 2023
     Experienced 0.62 0.49 0 1 1
     Pkg. contributions 78.06 563.84 0 0 19514
     Pkg. owner 0.07 0.26 0 0 1
     Followers 49.65 300.24 0 12 17421
     Following 36.54 313.69 0 11 28300
     Repositories 40.40 78.31 0 23 3900
     Total events 2.89 7.47 0 1 98
  • Notes:
  • Panel A presents descriptive statistics for the baseline sample period Pre 27-30.03 – Post 03-06.04. The “Less experienced” sample includes all GitHub user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of unique GitHub user accounts for the entire repository $\times$ user sample (“Overall”) and each of the subsamples is presented in parentheses below. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Panel B provides information on the individual characteristics of all GitHub user accounts in the repository $\times$ user sample.
(1) (2) (3) (4) (5)
A: Output Quantity and Quality
Output Output (Shen, 2023) Output (Holub & Thies, 2023 Issue closed PR merge ratio
Overall Treated $\times$ Post -0.0010 -0.0007 0.0029* -0.0007 -0.0002
(N = 239,912) (0.0010) (0.0010) (0.0016) (0.0005) (0.0005)
Dep. var. mean 0.0116 0.0108 0.0279 0.0024 0.0030
Less experienced Treated $\times$ Post 0.0019 0.0014 0.0047** 0.0007 0.0014***
(N = 86,200) (0.0015) (0.0014) (0.0022) (0.0006) (0.0005)
Dep. var. mean 0.0080 0.0075 0.0192 0.0014 0.0013
Experienced Treated $\times$ Post -0.0027** -0.0019 0.0018 -0.0016** -0.0011
(N = 153,712) (0.0014) (0.0013) (0.0022) (0.0007) (0.0007)
Dep. var. mean 0.0136 0.0128 0.0327 0.0030 0.0039
Pkg. contributor Treated $\times$ Post -0.0036 -0.0021 0.0011 -0.0025** 0.0001
(N = 63,064) (0.0025) (0.0024) (0.0036) (0.0012) (0.0011)
Dep. var. mean 0.0171 0.0163 0.0405 0.0034 0.0047
B: Task Choice and Complexity
PR opened Avg. lines added per opened PR Avg. lines added per merged PR Easy issue closed Interactive Activity
Overall Treated $\times$ Post 0.0008 0.0022 -0.0008 0.0001 0.0027**
(N = 239,912) (0.0007) (0.0027) (0.0020) (0.0001) (0.0013)
Dep. var. mean 0.0054 0.0170 0.0104 0.0001 0.0193
Less experienced Treated $\times$ Post 0.0019* 0.0059 0.0073*** 0.0000 0.0028
(N = 86,200) (0.0011) (0.0044) (0.0024) (0.0000) (0.0018)
Dep. var. mean 0.0041 0.0134 0.0047 0.0000 0.0127
Experienced Treated $\times$ Post 0.0001 0.0002 -0.0058** 0.0001 0.0026
(N = 153,712) (0.0010) (0.0034) (0.0029) (0.0002) (0.0018)
Dep. var. mean 0.0061 0.0190 0.0135 0.0001 0.0230
Pkg. contributor Treated $\times$ Post 0.0003 -0.0030 -0.0015 -0.0002 0.0031
(N = 63,064) (0.0016) (0.0058) (0.0043) (0.0002) (0.0031)
Dep. var. mean 0.0080 0.0238 0.0153 0.0001 0.0285
  • Notes:
  • All specifications include repository $\times$ user fixed effects and day-of-the-week–fixed effects. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Robust standard errors in parentheses are clustered on the repository $times$ user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
(1) (2) (3) (4) (5)
Output Output (Shen, 2023) Output (Holub & Thies, 2023 Issue closed PR merge ratio
Treated $\times$ Post -0.0042** -0.0022 -0.0007 -0.0021** -0.0011
(0.0021) (0.0021) (0.0040) (0.0009) (0.0007)
Overall (N = 33,512) Dep. var. mean 0.0078 0.0073 0.0227 0.0016 0.0016
Treated $\times$ Post -0.0004 0.0017 0.0027 -0.0019* 0.0004
(0.0029) (0.0028) (0.0057) (0.0010) (0.0007)
Less experienced (N = 12,784) Dep. var. mean 0.0055 0.0052 0.0173 0.0007 0.0003
Treated $\times$ Post -0.0069** -0.0048* -0.0031 -0.0023* -0.0021*
(0.0029) (0.0029) (0.0056) (0.0013) (0.0011)
Experienced (N = 20,728) Dep. var. mean 0.0093 0.0086 0.0261 0.0022 0.0023
Treated $\times$ Post -0.0060 -0.0024 0.0057 -0.0037** -0.0025
(0.0042) (0.0042) (0.0082) (0.0019) (0.0017)
Pkg. contributor (N = 11,392) Dep. var. mean 0.0121 0.0123 0.0364 0.0011 0.0020
  • Notes:
  • All specifications include repository $\times$ user fixed effects and day-of-the-week–fixed effects. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Robust standard errors in parentheses are clustered on the repository $times$ user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
Appendix E User Adaption to the ChatGPT Ban
Considering that there appears to be mean reversion in the estimated effects toward the end of our sample period, we now turn our attention to adaptation behaviour. The simplest way to circumvent the ChatGPT ban was to use VPN tools or encrypted routing through, for instance, the TOR network.
E.1 Data
[…]
Notes: The panel structure of the datasets used in Section LABEL:sec:google-tor—i.e., the (i) Google trends and the (ii) TOR (“standard” and bridge relay) user datasets—are displayed. Workday dates during the sample period are presented on the x-axis, and ISO 3166-1 alpha-2 country codes are presented on the y-axis.
E.2 Results
[…]
The top panel in Figure LABEL:fig:ban-circumvention presents the effect of the ChatGPT ban on the number of Google queries on the topic of VPNs. We observe a significant increase in the share of web searches on this relative to other topics in Italy on the first working day after the ban that slowly vanishes over the next three days. The estimated effect on April 3 is sizeable: the share of searches on VPNs increases by 52.2 percentage points. On average, the share of queries on VPNs was 20.6 percentage points higher in Italy over the workweek. The observed pattern is consistent with Italian users looking for ways to access ChatGPT even after the ban and succeeding after some initial search costs. Our estimates might, however, present only stated preferences.
[…]
To investigate whether the ban actually led to behavioral changes among Italian users, we look at an alternative outcome: the log number of TOR users. The results for TOR relay and TOR bridge users are presented in Panels B and C of Figure LABEL:fig:ban-circumvention, respectively. While the number of TOR relay users shows only a minor increase in the days after the ban, the average treatment effect on the number of Italian TOR bridge users is positive and significant on the first workday after the ban. Usage of TOR bridges remained elevated for the entire workweek, with an increase in user numbers—on average—of approximately 9.4 percentage points. This pattern is in line with users resorting to bridge over “standard” relays to minimize the chance of their being denied access to ChatGPT since the former are more difficult for firewalls to identify. [^36] For a discussion on denial of ChatGPT access, see the following OpenAI forum discussion: https://community.openai.com/t/access-denied-error-1020/38758/23.
Overall, our findings are consistent with Italian users looking for and finding ways to circumvent the blocked access to ChatGPT.
Table B.1: Variable Definitions (continued)
[…]